OpenStudy (anonymous):

Find polynomial p(x). Assume lead coefficient is one Degree 4, x=-1,x=2,x= i I worked it out but did not get the correct answer @precal

OpenStudy (jwheele1):

f(x) = a(x-c1)(x-c2)(x-c3)(x-c4)

OpenStudy (anonymous):

What you're given are the four roots of the function. We just need to rewrite them in factor forms.

OpenStudy (anonymous):

For instance, x=-1 in factor form is (x-(-1))=(x+1).

OpenStudy (anonymous):

x=2 would be (x-2)

OpenStudy (anonymous):

But then the x=i always come in pair. (x-i) (x+i)

OpenStudy (anonymous):

Putting them together, we should have... P=(x+1)(x-2)(x+i)(x-i)

OpenStudy (jwheele1):

Just now realized a = 1

OpenStudy (anonymous):

So I got (X+1)(x+2)(x^2+1)

OpenStudy (anonymous):

Oh dang I just seen what I did ... I got the sign wrong on the 2 Thank you very much for your help.

OpenStudy (anonymous):